OpenStudy (anonymous):

HELP PLEEEASEEEE!!!?? A rectangular pen with a perimeter of 60FT is needed to house three pigs. Determine an equation to model the area of this pen

OpenStudy (anonymous):

so the rectangle has length L and width W, if the perimeter is 60FT, this means that (L + W)*2 = 60, and Area = L x W

OpenStudy (anonymous):

2x+2y=60 x+y=30 x=30-y A=x*y A=(30-y)y

OpenStudy (anonymous):

as camp said 2(L+W)=60 because the perimeter is 60 so: L+W=30 Now the area is A=L*W so solve for 1 of ur dimensions in the perimeter equation: W=30-L or L=30-w and substitue into the area equation: A=(30-L)L or A=(30-W)W

OpenStudy (anonymous):

I still kinda confused :(

OpenStudy (anonymous):

ok u know that perimeter is adding up all the sides right. well there is 2 lengths and 2 widths. lets call the length L and the Width W so 2L+2W=P we know that P=60 so 2L+2W=60 so: L+W=30

OpenStudy (across):

Let L and W stand for length and width, respectively. What is the formula for the perimeter of a rectangle? \[P=2(L+W)\]It can't get any easier than that. What is the formula for the area of a rectangle? \[A=LW\]It can't get any easier than that. From the question, we know that: \[2(L+W)=60\]We can solve this for any of the two variables. For the sake of whatever, let's pick L: \[L=30-W\]This is basic algebra. Now we plug this new L into the area formula: \[A=W(30-W)\]And there you have it: The area of the pen in terms of its width.

OpenStudy (anonymous):

so the length added to the width equals 30

OpenStudy (anonymous):

Ok, so basically you're limited by the perimeter. So think about it this way, you have 60FT of rope to make a rectangular box with. (your pen for the pigs) Let's say you make the length of the box 20 FT, the width must be 10 FT, because the perimeter must add up to 60FT, from 2L + 2W = 2*20 + 2*10 = 60FT This means that the area is L*W = 20*10 = 200 FT squared

OpenStudy (anonymous):

I think i understand it now based on what across and camp wrote

OpenStudy (anonymous):

thanks for you help